# CapSpeech: Enabling Downstream Applications in Style-Captioned Text-to-Speech

### v1

@@ 3.2.2 SFT Sets for CapTTS, EmoCapTTS, and AccCapTTS
We group CapTTS, EmoCapTTS, and AccCapTTS under the same subsection because they share both underlying speech corpora and common stylistic attributes (e.g., timbre, speaking rate). To support these tasks, we aggregate human-annotated data from six publicly available corpora: LibriTTS-R DBLP:conf/interspeech/KoizumiZKDYMB0H23 , VCTK Yamagishi2019VCTK , VoxCeleb DBLP:conf/interspeech/NagraniCZ17 , VoxCeleb2 DBLP:conf/interspeech/ChungNZ18 , EARS DBLP:journals/corr/abs-2406-06185 , and Expresso DBLP:conf/interspeech/NguyenHDSGFRCSH23 . The aggregation process involves directly using existing captions, enhancing captions with additional speaker traits, and generating captions using a large language model based on structured labels. For LibriTTS-R, we incorporate annotations from LibriTTS-P DBLP:journals/corr/abs-2406-07969 and DreamVoiceDB hai2024dreamvoice , which particularly focus on intrinsic speaker traits, and use the LLM to generate captions. From DreamVoiceDB, we also use the VCTK portion, which includes accent annotations, and use the LLM to generate captions. Please refer to Appendix B.1 for the LLM prompts used in this process. For VoxCeleb, VoxCeleb2, EARS, and Expresso, we apply annotations from ParaSpeechCaps DBLP:journals/corr/abs-2503-04713 , which cover both intrinsic and expressive style traits.
@@ 3.2.4 SFT Set for AgentTTS
To collect the recordings, we first generate prompt-content pairs, paying careful attention to balancing subcategories of emotions and styles. The content spans a wide range of conversational scenarios, including everyday dialogue, cafés, hospitals, classrooms, sci-fi films, horror movies, and more. During recording, the speaker follows the prompts to perform the specified emotion and style, and is allowed to revise the prompt or content if any part is unclear or unnatural. The prompts used could be found in Appendix B.3.
@@
4 Experiments
4.1 Baselines
We adopt two strong baselines for the CapSpeech benchmark, which is shown in Figure 2. See Appendix C for the model complexity.
@@ 4.1 Baselines
CapSpeech-AR: The first is Parler-TTS [^4] https://github.com/huggingface/parler-tts DBLP:journals/corr/abs-2402-01912 , a state-of-the-art AR method based on a codec language model. We use the 44.1kHz version of Descript Audio Codec (DAC) DBLP:conf/nips/KumarSLKK23 to provide discrete audio representations. A delay pattern is applied to handle multiple codebooks DBLP:conf/nips/CopetKGRKSAD23 , and cross-attention is used to incorporate caption-based style conditioning. FLAN-T5555https://huggingface.co/google/flan-t5-large DBLP:journals/jmlr/ChungHLZTFL00BW24 is used to extract the text features from the transcription and style caption. We retain the original architecture, with the exception of adding special tokens to support the CapTTS-SE task. The FLAN-T5 model used in the transcription is fine-tuned. As illustrated in Figure 1, transcriptions in CapTTS-SE support two modes for integrating sound events: (i) Background mode and (ii) Insertion mode. sound event tokens (e.g., <telephone>, <knock>) are placed at the beginning of the sequence. The tags <B> and </B> mark the start and end of a background sound segment, while <I> </I> denote the insertion point. This design allows flexible control over the placement and type of sound events within the synthesized speech.
@@
CapSpeech-NAR: In addition, we adopt F5-TTS [^6] https://github.com/SWivid/F5-TTS DBLP:journals/corr/abs-2410-06885 , a state-of-the-art NAR method based on flow matching with a Diffusion Transformer. In our adaptation, we remove the audio-prompt masking component and instead use cross-attention to incorporate caption-based style conditioning. BigVGAN [^7] https://huggingface.co/nvidia/bigvgan_v2_24khz_100band_256x DBLP:conf/iclr/LeePGCY23 is used as the vocoder, and we apply QK-Norm DBLP:conf/emnlp/HenryDPC20 to stabilize training. Transcriptions are processed using grapheme-to-phoneme conversion (g2p) [^8] https://github.com/Kyubyong/g2p, and we insert special tokens <B>, </B>, <I>, and </I> to indicate background and insertion-based sound events. To enhance generalization, we do not include sound event tags directly in the input sequence. Instead, we extract LAION-CLAP DBLP:conf/icassp/WuCZHBD23 embeddings for the specified sound event and use them as the additional input. This design allows the model to generalize to unseen sound events during inference. Similar to the AR model, a Flan-T5 model is used to extract the text features from the style caption. Since the NAR model cannot predict audio duration directly, inspired by DBLP:journals/corr/abs-2408-13893 , we fine-tune a BERT model [^9] https://huggingface.co/google-bert/bert-base-uncased DBLP:journals/corr/abs-1810-04805 that takes both the transcription and the caption as input to estimate the total duration of the entire audio.
4.2 Implementation Details
@@
All models are trained using the AdamW optimizer DBLP:conf/iclr/LoshchilovH19 . The AR model is trained with a batch size of 32 and a learning rate of 1e-3 during pretraining, and 1e-4 during fine-tuning. The NAR model uses a batch size of 512, with a learning rate of 2e-4 for pretraining and 2e-5 for fine-tuning. pretraining is conducted on 8 NVIDIA H100 GPUs, while fine-tuning is performed on a single NVIDIA A100 GPU. Following DBLP:journals/corr/abs-2503-04713 , we initialize the AR model with the Parler-TTS Mini v1101010https://huggingface.co/parler-tts/parler-tts-mini-v1 model. During inference, we perform AR generation with a temperature of 1.0 and a repetition penalty of 1.0, and introduce inference-only classifier-free guidance (CFG) DBLP:journals/corr/abs-2409-07556 . For the NAR model, following DBLP:journals/corr/abs-2410-06885 , we use the default CFG strength of 2 and a Sway Sampling coefficient of $-1$. Please refer to the Appendix D for additional implementation details.
4.3 Metrics
@@
Objective Metrics: We evaluate three key aspects of model performance: style consistency, audio quality, and intelligibility. For style consistency, we compute classification accuracy across multiple categories: age, gender, pitch, expressiveness of tone, speed, accent, and emotion. The average accuracy across these attributes is reported as Style-ACC. Please refer to the Appendix F for detailed procedures. Audio quality is assessed using UTMOSv2111111https://github.com/sarulab-speech/UTMOSv2 DBLP:conf/slt/BabaNSS24 . For intelligibility, we compute a text-normalized WER between the ASR transcript of the generated speech and the input transcript. This is done using openai/whisper-large-v3-turbo along with the Whisper text normalizer.
Subjective Metrics: We experimented with state-of-the-art speech understanding models DBLP:journals/corr/abs-2407-10759 ; kimiteam2025kimiaudiotechnicalreport to evaluate style consistency, but they failed to generate high-quality speech captions. Following DBLP:journals/corr/abs-2503-04713 , we recruit 15 native speakers via Prolific [^12] https://www.prolific.com/ to evaluate three subjective aspects: Style Consistency MOS (SMOS), Naturalness MOS (NMOS), and Intelligibility MOS (IMOS). Each example is rated by three annotators, and we report the mean scores along with $95\%$ confidence intervals. See Appendix G for details.
4.4 Results and Analysis
Table 3: Results on the pretraining task.
@@ 4.4 Results and Analysis
[…]
NAR ParaSpeechCaps $51.8\%$ $3.16$ $9.5\%$ $3.55{\scriptstyle\pm 0.11}$ $3.38{\scriptstyle\pm 0.09}$ $3.93{\scriptstyle\pm 0.10}$
[…]
Pretraining Stage: We compare the CapTTS task trained on our proposed CapTTS pretraining set against training on the previous large-scale dataset ParaSpeechCaps (PSC-Scaled partition). Table 3 presents the results on the CapTTS task, comparing models using both objective and subjective metrics. Details of the Style-ACC scores can be found in Appendix E. Models trained on the CapTTS Pretraining set achieve significantly better style consistency, naturalness, and intelligibility than those trained on ParaSpeechCaps, demonstrating the effectiveness of our dataset. Compared to AR models, NAR models consistently achieve better performance across all metrics, highlighting their advancement on the CapTTS task.
Fine-tuning Stage: We pre-train the AR and NAR models on our proposed CapTTS and CapTTS-SE pretraining sets, and then fine-tune them on the downstream tasks. Details of the Style-ACC scores can be found in Appendix E. The results, shown in Table 4, demonstrate that pretraining provides substantial benefits across all downstream tasks–particularly for CapTTS-SE and AgentTTS, where training data is limited. Notably, our benchmark indicates that strong style consistency, naturalness, and intelligibility can be achieved on the CapTTS, EmoCapTTS, and AccCapTTS tasks, with the NAR model reaching SMOS, NMOS, and IMOS scores of at least 3.77, 3.88, and 4.34, respectively. Additionally, the AR model surpasses the NAR model on certain metrics in the CapTTS-SE and AgentTTS tasks. We observe that maintaining style consistency in the AgentTTS task and achieving high intelligibility in the CapTTS-SE task remain particularly challenging. In particular, models achieve good WER results but perform poorly on IMOS in the CapTTS-SE task, indicating that sound events are generated with lower quality than speech.
Table 4: Results on the fine-tuning tasks. $\star$ denotes models without pretraining. CapTTS-SE results without pretraining are not reported as there are only 500 samples in CapTTS-SEDB. Style-ACC is not applicable to AgentTTS and CapTTS-SE, and UTMOSv2 is not applicable to CapTTS-SE.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
5 Conclusion and Discussions

### v2

@@ III-B2 SFT Sets for CapTTS, EmoCapTTS, and AccCapTTS
We group CapTTS, EmoCapTTS, and AccCapTTS together because they share both underlying speech corpora and common stylistic attributes (e.g., timbre, speaking rate). To support these tasks, we aggregate human-annotated data from six publicly available corpora: LibriTTS-R [32], VCTK [36], VoxCeleb [33], VoxCeleb2 [34], EARS [35], and Expresso [20]. The aggregation process involves directly using existing captions, enhancing captions with additional speaker traits, and generating captions using a large language model based on structured labels. For LibriTTS-R, we incorporate annotations from LibriTTS-P [26] and DreamVoiceDB [42], which particularly focus on intrinsic speaker traits, and use the LLM to generate captions. The prompt used for generating LibriTTS-P annotations is shown below.
Prompt Design for LibriTTS-P System: You are a voice style caption generator. User: • Given a list of keywords related to a speaker’s vocal traits, generate 10 diverse and natural-sounding style captions. • You can freely drop or combine keywords. Vary the sentence structures and expressions as much as possible. • Respond with only the caption. Keywords: [e.g., teenager, female, bright, smooth, nasal, cute, and quick pace.]
LLM Assistant: [e.g., A teenage girl’s voice is bright and smooth, with a slight nasal quality, and she speaks at a lively, quick pace.]
From DreamVoiceDB, we also use the VCTK portion, which includes accent annotations, and use the LLM to generate captions. The prompt used for this caption generation is shown below.
Prompt Design for VCTK System: You are a voice style caption generator. User: • Please naturally incorporate the accent information into the following description, without changing its original meaning. Accent: [e.g., Indian Accent.] Original description: [e.g., A senior woman’s voice carries with warmth, depth, and an authoritative tone.] • You can freely vary the sentence structures and expressions as much as possible. • Respond with only the modified caption. LLM Assistant: [e.g., A senior woman’s voice resonates with warmth and depth, layered with an authoritative tone and gently marked by an Indian accent.]
For VoxCeleb, VoxCeleb2, EARS, and Expresso, we apply annotations from ParaSpeechCaps [25], which cover both intrinsic and expressive style traits.
@@ III-B4 SFT Set for AgentTTS
To collect the recordings, we first generate prompt-content pairs, paying careful attention to balancing subcategories of emotions and styles. The content spans a wide range of conversational scenarios, including everyday dialogue, cafés, hospitals, classrooms, sci-fi films, horror movies, and more. During recording, the speaker follows the prompts to perform the specified emotion and style, and is allowed to revise the prompt or content if any part is unclear or unnatural.
The following prompt template guides an LLM in generating expressive prosody descriptions and corresponding dialogue samples. As this task is more complex, we use GPT-4o-mini [^8] https://platform.openai.com/docs/models/gpt-4o-mini for better quality and naturalness. Below is the prompts used in this process.
Prompt Design for CapSpeech-AgentDB System: You are a speech and voice expert contributing to an expressive speech dataset. User: Please generate a natural-language prosody prompt that vividly describes how the speaker’s voice would sound, given a high-level expression (e.g., emotion, mental state) and a low-level expression (e.g., pitch, pace, volume). Guidelines: • Ensure that the prosody description is vivid, coherent, and captures the expression naturally. • It is acceptable to merge or omit specific low-level features for fluency. • Please follow these examples: [random sample 3 human-written style-captions]
LLM Assistant: [e.g., The voice is quick and unsteady, occasionally faltering, revealing anxiousness.]
User: Using the above prosody prompt and the provided conversation scene, generate a single line of dialogue or a short dialogue excerpt from only one speaker.
Guidelines: • Ensure the speech sounds natural and is emotionally aligned with the described vocal tone and context. • Use varied sentence structures and lexical choices across examples to enhance dataset diversity. • Do not include or imply any lines from other characters—generate content exclusively for one speaker. Optional: When appropriate, incorporate non-speech vocalizations from the provided list ([Gasp], [Laughter], …, [Yawn]) to enhance the expressiveness of the speech. Conversation scene: [e.g., in a horror movie]
LLM Assistant: [e.g., I—I think I saw something move behind the curtain. No, no, I’m not imagining it. It was real. I swear it was real. We can’t stay here. We have to go—now!]
These outputs form the basis of the CapSpeech-AgentDB, providing reference material for voice actors to produce the final audio recordings. The set of high-level emotional expressions, low-level prosody expressions, and non-speech vocalizations used to guide the prompts is presented in Table V.
TABLE V: List of Expressive Voice Labels
High-level Emotional Expressions Neutral Aggressive Angry Annoyed Contemptuous Hateful Raging
Disgusted Sarcastic Resentful Alarmed Apprehensive Fearful Nervous
Panicked Terrified Accepting Admiring Amazed Excited Hopeful
Interested Joyful Loving Peaceful Surprised Disappointed Dismayed
Distressed Grieving Hopeless Sad Authoritative Respectful Sympathetic
Trusting Anticipating Curious Eager Focused Tired Awkward
Confused Distracted Skeptical Bored Envious Submissive Shameful
Low-level Prosody Expressions Default Monotonous Deep Sharp Gentle Loud Mumble
Stutter Whispering Crying Laughing
Non-speech Voices Aww Throat-clearing Cheering Contemplation Gasp Groan Laughter
Panting Scream Sigh Sneering laughter Sob Yawn
@@
IV Experiments
IV-A Baselines
We adopt two strong baselines for the CapSpeech benchmark.
@@ IV-A Baselines
CapSpeech-AR: The first is Parler-TTS [^9] https://github.com/huggingface/parler-tts [23], a state-of-the-art AR method based on a codec language model. We use the 44.1kHz version of Descript Audio Codec (DAC) [58] to provide discrete audio representations. A delay pattern is applied to handle multiple codebooks [59], and cross-attention is used to incorporate caption-based style conditioning. FLAN-T5101010https://huggingface.co/google/flan-t5-large [60] is used to extract the text features from the transcription and style caption. We retain the original architecture, with the exception of adding special tokens to support the CapTTS-SE task. The FLAN-T5 model used in the transcription is fine-tuned. As illustrated in Fig. 1, transcriptions in CapTTS-SE support two modes for integrating sound events: (i) Background mode and (ii) Insertion mode. sound event tokens (e.g., <telephone>, <knock>) are placed at the beginning of the sequence. The tags <B> and </B> mark the start and end of a background sound segment, while <I> </I> denote the insertion point. This design allows flexible control over the placement and type of sound events within the synthesized speech.
@@
CapSpeech-NAR: In addition, we adopt F5-TTS [^11] https://github.com/SWivid/F5-TTS [8], a state-of-the-art NAR method based on flow matching with a Diffusion Transformer. In our adaptation, we remove the audio-prompt masking component and instead use cross-attention to incorporate caption-based style conditioning. BigVGAN [^12] https://huggingface.co/nvidia/bigvgan_v2_24khz_100band_256x [61] is used as the vocoder, and we apply QK-Norm [62] to stabilize training. Transcriptions are processed using grapheme-to-phoneme conversion (g2p) [^13] https://github.com/Kyubyong/g2p, and we insert special tokens <B>, </B>, <I>, and </I> to indicate background and insertion-based sound events. To enhance generalization, we do not include sound event tags directly in the input sequence. Instead, we extract LAION-CLAP [63] embeddings for the specified sound event and use them as the additional input. This design allows the model to generalize to unseen sound events during inference. Similar to the AR model, a Flan-T5 model is used to extract the text features from the style caption. Since the NAR model cannot predict audio duration directly, inspired by [4], we fine-tune a BERT model [^14] https://huggingface.co/google-bert/bert-base-uncased [64] that takes both the transcription and the caption as input to estimate the total duration of the entire audio.
IV-B Implementation Details
Table VI summarizes the computation cost and model size of AR and NAR models. In the Transformer column, the numbers denote the model dimension, the number of layers, the number of heads, and the multiples of hidden size. Training time is calculated on 8 NVIDIA-H100 GPUs and Real-time-factor (RTF) is calculated on a single NVIDIA-A100 GPU.
TABLE VI: Model complexity and computational cost.
Model Transformer Params(M) Training Time RTF
CapSpeech-AR 1024,24,16,4 880 24 Days 1.24
CapSpeech-NAR 1024,24,16,4 724 10 Days 1.01
@@ IV-B Implementation Details
All models are trained using the AdamW optimizer [65]. Following [25], we initialize the AR model with the Parler-TTS Mini v1151515https://huggingface.co/parler-tts/parler-tts-mini-v1 model. During inference, we perform AR generation with a temperature of 1.0 and a repetition penalty of 1.0, and introduce inference-only classifier-free guidance (CFG) [66]. For the NAR model, following [8], we use the default CFG strength of 2 and a Sway Sampling coefficient of 1-1. To avoid any potential misuse, we introduce a novel watermarking technology [66] to each model, and open-source our code and model weights under CC BY-NC 4.0 license.
IV-C Metrics
Objective Metrics: We evaluate three key aspects of model performance: style consistency, audio quality, and intelligibility. For style consistency, we compute classification accuracy across multiple categories: age, gender, pitch, expressiveness of tone, speed, accent, and emotion. The average accuracy across these attributes is reported as Style-ACC. Following Section III-B1, we utilized existing toolkits to predict various speaker attributes, including age, gender, pitch, vocal expressiveness, and speaking rate. For emotion and accent classification, we introduced the state-of-the-art classifiers in [67]. The emotion classifier includes nine emotion labels: Anger, Contempt, Disgust, Fear, Happiness, Neutral, Sadness, Surprise, and Other. The accent classifier covers 16 accent categories: East Asia, English, Germanic (e.g., German), Irish, North America, Northern Irish, Oceania (e.g., Australia), Romance (e.g., Spanish, French), Scottish, Semitic, Slavic (e.g., Russian, Polish), South African, Southeast Asia, South Asia, Welsh, and Other accents not covered above.
@@ IV-C Metrics
Audio quality is assessed using UTMOSv2161616https://github.com/sarulab-speech/UTMOSv2 [68]. For intelligibility, we compute a text-normalized WER between the ASR transcript of the generated speech and the input transcript. This is done using openai/whisper-large-v3-turbo along with the Whisper text normalizer.
[…]
IV-D Results and Analysis
TABLE VII: Results on the pretraining task.
@@ IV-D Results and Analysis
Ours $\boldsymbol{52.2\%}$ $\boldsymbol{3.18}$ $\boldsymbol{9.1\%}$ $\boldsymbol{3.62}{\scriptstyle\pm 0.11}$ $\boldsymbol{3.65}{\scriptstyle\pm 0.10}$ $\boldsymbol{4.23}{\scriptstyle\pm 0.12}$
NAR ParaSpeechCaps $51.8\%$ $3.16$ $9.5\%$ $55{\scriptstyle\pm 0.11} 3.3$$38{\scriptstyle\pm 0.09}$ $3.93{\scriptstyle\pm 0.10}$
Ours $\boldsymbol{62.1\%}$ $\boldsymbol{3.43}$ $\boldsymbol{8.8\%}$ $\boldsymbol{3.80}{\scriptstyle\pm 0.13}$ $\boldsymbol{3.79}{\scriptstyle\pm 0.10}$ $\boldsymbol{4.40}{\scriptstyle\pm 0.09}$
TABLE VIII: Results on the fine-tuning tasks. $\star$ denotes models without pretraining. CapTTS-SE results without pretraining are not reported as there are only 500 samples in CapTTS-SEDB. Style-ACC is not applicable to AgentTTS and CapTTS-SE, and UTMOSv2 is not applicable to CapTTS-SE.
@@
NAR $\boldsymbol{66.0\%}$ $\boldsymbol{3.37}$ $\boldsymbol{9.2\%}$ $\boldsymbol{3.85\scriptstyle\pm 0.13}$ $\boldsymbol{3.95\scriptstyle\pm 0.12}$ $\boldsymbol{4.34\scriptstyle\pm 0.11}$
@@
NAR / / $\boldsymbol{3.0\%}$ $\boldsymbol{3.75\scriptstyle\pm 0.13}$ $\boldsymbol{3.60\scriptstyle\pm 0.11}$ $3.33\scriptstyle\pm 0.15$
@@
NAR $\boldsymbol{67.2\%}$ $\boldsymbol{3.34}$ $\boldsymbol{10.1\%}$ $\boldsymbol{3.77\scriptstyle\pm 0.12}$ $\boldsymbol{3.88\scriptstyle\pm 0.13}$ $\boldsymbol{4.35\scriptstyle\pm 0.10}$
@@
NAR $\boldsymbol{66.4\%}$ $\boldsymbol{3.45}$ $\boldsymbol{8.8\%}$ $\boldsymbol{3.91\scriptstyle\pm 0.12}$ $\boldsymbol{3.90\scriptstyle\pm 0.12}$ $\boldsymbol{4.45\scriptstyle\pm 0.09}$
@@
NAR / $3.07$ $\boldsymbol{9.5\%}$ $3.42\scriptstyle\pm 0.14$ $\boldsymbol{3.80\scriptstyle\pm 0.11}$ $\boldsymbol{4.41\scriptstyle\pm 0.10}$
Pretraining Stage: We compare the CapTTS task trained on our proposed CapTTS pretraining set against training on the previous large-scale dataset ParaSpeechCaps (PSC-Scaled partition). Table VII presents the results on the CapTTS task, comparing models using both objective and subjective metrics. Models trained on the CapTTS Pretraining set achieve significantly better style consistency, naturalness, and intelligibility than those trained on ParaSpeechCaps, demonstrating the effectiveness of our dataset. Compared to AR models, NAR models consistently achieve better performance across all metrics, highlighting their advancement on the CapTTS task.
Fine-tuning Stage: We pre-train the AR and NAR models on our proposed CapTTS and CapTTS-SE pretraining sets, and then fine-tune them on the downstream tasks. The results, shown in Table VIII, demonstrate that pretraining provides substantial benefits across all downstream tasks–particularly for CapTTS-SE and AgentTTS, where training data is limited. Notably, our benchmark indicates that strong style consistency, naturalness, and intelligibility can be achieved on the CapTTS, EmoCapTTS, and AccCapTTS tasks, with the NAR model reaching SMOS, NMOS, and IMOS scores of at least 3.77, 3.88, and 4.34, respectively. Additionally, the AR model surpasses the NAR model on certain metrics in the CapTTS-SE and AgentTTS tasks. We observe that maintaining style consistency in the AgentTTS task and achieving high intelligibility in the CapTTS-SE task remain particularly challenging. In particular, models achieve good WER results but perform poorly on IMOS in the CapTTS-SE task, indicating that sound events are generated with lower quality than speech.
V Conclusion and Discussions